OpenStudy (anonymous):

What is the sum of the geometric sequence –4, 24, –144, ... if there are 6 terms? How do I do this?

OpenStudy (meehan98):

Ok, the equation looks like this: \[S _{n}=a _{1} \frac{ 1-r ^{n} }{ 1-r }\]

OpenStudy (anonymous):

what's S, A, R, and N?

OpenStudy (meehan98):

S= the sum n=the term that you choose, so for this example it would be 6 \[a _{1}\]= the first term

OpenStudy (anonymous):

So exactly how would you set it up? thats all I need to know, can you show me an example?

OpenStudy (meehan98):

Let's break it down a little bit, so it's not so much to take in. We just fill in what we know so: \[S _{6}= -4 \frac{ 1-(-6)^{6} }{ 1-(-6)}\]

OpenStudy (meehan98):

I forgot to tell you that we get "r" by 24/-4=-6 because that's our ratio.

OpenStudy (drigobri):

This equation does not work @Meehan98 If you try to do for the second term (S2), will hit as the number 20, which is not consistent with the case

OpenStudy (anonymous):

Just find the pattern In this case its multiplying by -6

OpenStudy (anonymous):

Sohal i'm in trigonometry, i'm not just doing x-6 lol

OpenStudy (anonymous):

so the answer would be... 864,-5184,31104 would be the next numbers

OpenStudy (anonymous):

Lol but all the information is informant of you -_- and the way you get the answer doesn't matter as long as you don't cheat,show work,and get the right answer

OpenStudy (anonymous):

do you deny that I got the right answer?

OpenStudy (anonymous):

I found the answer deductivly

OpenStudy (anonymous):

Ps I never actually got the answer But I did give you the next numbers of the sequence

OpenStudy (meehan98):

hmm..this doesn't make sense then. I've been doing these problems all day without an issue..

OpenStudy (anonymous):

ok do it your way ill do it my way -_- honestly ever herd of doing it in different methods forget it

OpenStudy (meehan98):

I know; because the formula that I have up there is the Sum Formula so it wouldn't make sense to fit the second term in there. The formula for the terms is:\[a _{n}=a _{1} r ^{n-1}\]

OpenStudy (drigobri):

If you get the second number and divide by the first, will result 6. If you made the third number and divide by the first, will result in 36, which is 6 ^ 2, we can quickly realize the equation, that is: \[n = 4 x (-6^{(n-1)})\] where: n = number wish to find So, \[n1= -4 . (-6^{1-1}) = -4 . (1) = -4\]\[n2= -4 . (-6^{2-1}) = -4 . (-6) = 24\]\[n3= -4 . (-6^{3-1}) = -4 . (36) = -144\]\[n6= -4 . (-6^{6-1}) = -4 . (-46656) = +186624\]

OpenStudy (drigobri):

we can quickly realize the equation, that is: n=-4x(−6(n−1))***** << I forgot -4 there